OpenStudy (anonymous):

maximize F=x^2+y^2 with x + 2y = 10.

OpenStudy (anonymous):

Are you using Lagrange multipliers?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Okay well clearly: \[ x = 10-2y \]So \[ F(y) = (10-2y)^2+y^2 \]Find \(F'(y)\).